2021 August/September Restart OUTLINE

September 1-12 2021 - Engineering On-Sky Restart Block


General Scope of Activities

  • Closed Dome/Daytime Instrument Checks (each instrument must be released before night-time operations)
  • Telescope Wake-Up
  • Telescope Infrastructure
    • AGW1/2
    • M3 SX/DX
  • Checkout of Facility Instruments
  • LBTI Restart (4 x 0.5 nights) - pending SX ASM release
  • AO (SX) Restart (2 x 0.5 nights) -pending SX ASM release

Key Dates & Information

  • DONE SX M3 Install 8/12
  • DONE DX M3 Install 8/27
  • DONE AGW-1 Install 8/13
  • DONE AGW-2 Install 9/2
  • DONE LUCI-1 Install Complete 8/16
  • LBTI activities complete 8/26
  • DONE LBC-B Install Complete 8/30
  • DONE LBC-R Install Complete 8/30
  • DONE MODS-1 Install Complete 8/23
  • DONE MODS-2 Install Complete 8/30
  • DONE LUCI-2 Install Complete 9/20
  • LBTI cold 9/3
  • SX ASM AO not likely to be available 9/1-9/12 (SL will be available)
    • Earliest closed dome testing now estimated 10/14-10/15

  • RESTART NIGHTS:
    • Sep 1-Sep 7
    • OSURC Sep 8-9
    • Partial PETS/Restart Sep 10
      • PETS Twilight-06:43 UT (01:43 local time)
    • Sep 11-15

Personnel

  • Night Activities

  • Daytime Activities

Closed Dome/Daytime Work

  • LBC (2 hrs)
    • DONE Filter Functional checks
    • DONE Rotator Functional checks
    • DONE Test Bias Exposures
  • MODS
    • Functional Checkout (2hrs)
      • DONE M1
      • DONE M2
    • Test Exposures (sieve and internal flat snaps)
      • DONE M1
      • DONE M2
    • Internal Calibrations (2hrs)
      • M1
        • Grating Mode
        • Prism Mode
        • Imaging Mode
      • M2
        • DONE Grating Mode
        • Prism Mode
        • DONE Imaging Mode
    • MODS Mask Exchange
      • DONE M1
      • DONE M2
    • MODS Focus Checks
      • DONE M1
      • DONE M2 - all modes (ran UT Sep 02 and UT Sep 05!)
  • AGw1/AGw2 aliveness check (1 hr)
    • DONE take images with all cameras (could be closed-dome preset)
    • DONE move stages
  • LUCIs (5-6 days total)
    • Basic Checks needed for release for engineering (and AO) work - 1 day i n total
      • DONE non-MOS Functional Checkout on telescope L1 (cold) - 0.5 day
      • DONE non-MOS Functional Checkout on telescope L2 (cold) - 0.5 day
      • DONE MOS Functional Checkout on telescope L1 (cold) - 0.5 day
      • DONE MOS Functional Checkout on telescope L2 (cold) - 0.5 day
      • DONE PFS alignment and Camera and Grating refocus L1 - 1hr
        • may need to re-check G200
      • DONE PFS alignment and Camera and Grating refocus L2- 1hr
    • Grating Wavelength Calibration - 3-5hr take data, 1-2 days process data
      • DONE G210 L1
      • G210 L2
      • DONE G200 L1
      • G200 L2
      • G150 L1 (low priority)
      • G040 L2 (lowest priority)
    • Flexure Checks - 3 days
      • DONE / DONE - G210/N1.8 - L1/L2
      • G210/N3.75 - L1/L2
      • G200/N3.75 - L1/L2
      • DONE / DONE - G200/N1.8 - L1/L2
      • G150/N1.8 - L1
      • N1.8 imaging - L1/L2
      • DONE / TBD - N3.75 imaging - L1/L2
      • DONE / DONE N30 Imaging - L1/L2
    • Mask Exchange
      • DONE L1
      • DONE L2
  • LBTI (2 days)
    • Internal checkout (0.5 day)
    • AdSec check out (1.5 day)
  • AO (3-4 days)
    • AO flat modes checkout (to verify SL use) - 1 day
    • IF AO is to be used
      • 4D alignment (2 x 0.5 days or 1 day)
      • AO WFS checkout (internal)
      • AO flats - (2 x 0.5 sharaed with LUCI checkout)
      • Full AO checkout - (2 x 0.5 days - shared with LUCI checkout)
      • Full AO Checkout with LUCI - including NCPA (2 x 0.5 day - shared with LUCI Cabex's)
      • ARGOS cal unit checkout
  • PEPSI
    • Internal Operations (on-sky with solar telescope already)
    • Test presets with MODS/PEPSI mixed mode

List of On-Sky Tasks

  • Telescope Wake-up (1 hr)
    • Closed Dome Test Presets (multiple instruments)
      • DONE LBC
      • DONE LUCIs
    • On-sky presets (Track, Acquire, Active)
      • DONE LBC Track
      • DONE LUCIs Active, Acquire
      • DONE MODS Active, Acquire
  • Telecope Infrastructure
    • DONE SX M3 Alignment LFBG - Critical (3 hrs) - requires a LUCI AGw - both sides can be done in parallel
      • Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout, but active optics needs to be working
      • Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
    • DONE DX M3 Alignment RFBG - Critical (3 hrs) - requires a LUCI AGw
      • Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout
      • Pupil wobble (2 hrs) - requires LUCI-2/AGW-2 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
    • DONE SX M3 Alignment PFU - check
    • DONE DX M3 Alignment PFU - check
    • Field Aberration (2 hrs) - Medium Priority (1hour) (DX more important than SX)
      • DONE DONE SX LUCI data collected
      • DONE DX LUCI data collected
      • DONE SX MODS data collected
      • DONE DX MODS data collected
        • Take test data if conditons are poor to verify scripts - Can be done with MODS early in restart
        • Might be better to use LUCIs due to IT 8315
          • Just need both AGW1 and AGW2 on the telescope and checked to execute
        • requires AGW-1/AGW-2 or AGW5/6 to guide, collimate, SH spots well illuminated, hotspot centered - can be done before AGW Rot ZP and Transform
        • IDL routines fixed
        • Requires < 1" to get good results
          • Don't hold out for 1", take in 1.2" to get first order results.
      • DONE SX data analyzed (DLM 20210908)
      • DONE DX data analyzed
      • DONE SX correction installed
      • DONE DX correction installed - Not needed - no changes
      • DONE SX LUCI check data collected
      • DONE DX LUCI check data collected
  • LBC (high)
    • DONE Collimate WRS/DOFPIA - (4 x 0.25 hrs)
      • DONE wouldn't hurt to collimate in better conditions
    • DONE Superfocus sequence -PA=0,180, 90, 270 - high (1 hr)
    • DONE Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr)
      • two airmasses (1.64,1.04)
    • DONE Test Science OB - guiding - medium (1 hr)
      • DONE Sidereal
        • DONE cycled through 3 filters in each LBC and dithered once - HCG 98wouldn't hurt to complete an OB
        • DONE HCG 98 completed all but last dither position
      • DONE Non-Sidereal OB - Comet 67P/Churyumov-Gerasimenko
    • DONE Standard photometric fields (new ones ready) - requires clear skies (any seeing) - required for ZP calibrations
      • UT Sep 15 - 3 sets
      • UT Sep 16 - 1 set
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • DONE Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
    • on-sky rotational trailing tests
      • To be done AFTER a new pointing model implemented
    • TMS with LBC (Heejoo, Olga and John H.)
      • DONE collect passive TMS data during LBC observations medium for all LBC observations
      • DONE use active TMS for some 1 hour+ LBC observations medium/high
      • use active TMS plus Andrew's thermal correction for some 1 hour+ LBC observations (still needs a script to apply the thermal correction)
    • DONE LBC Pointing Model - (3hr) - any conditions
      • DONE TMS (active) - low/medium
        • unsure if pointing model can be generated if run in active mode
        • timing problems - slewing vs TMS inputing corrections (JMH may have fixed this with his update to lbc_tms.py on 20210905), so we should at least try the active TMS pointing model
      • DONE Standard method -medium/High
        • PtStarrs - please choose targets > 80 deg elevation (more than 1 or 2)
        • Run TMS passively - run 1 reference at the beginning
        • Need stable temperatures
        • Turn on PCS logging
        • Need to collimate at each star
        • Chip 2 only exposure
        • Analysis done by - O. Kuhn
        • File Location:
          • /home/lbto/Calib_OBs/PtStars_10m_OBs/VV

    • Cryostat tilt adjustment check
      • Use the superfoc data
      • Calculate shim thickness from data - off telescope
      • Iterations might be needed - 1hr per iteration
  • MODS (high)
    • DONE Collimate AGWs
    • DONE Check SH spots
    • DONE check MODS1/MODS2 guide probe wrt/rotator center
      • Files needed:
        • agw.ini located on mods1 and mods2 computers in: /home/mods/Config/MODS1 or /home/mods/Config/MODS2
    • DONE MODS Rotator Center measure (use M71 at beg of night or M52 later)
      • requires pointing mask (40" square field stop & dot-circle with radius 45") - high (0.5hr)
    • DONE MODS AGW 5/6 - Measure Rotator zero point (mask alignment) - high (0.5 hr) - mask exchange required
      • M1: dPA = mean theta = -0.1491 (-0.1617,-0.1282, -0.1574) -updated PCSInstrument.conf for Transform collection
        • left rzp from 27.63 to 27.78
      • M2: dPA = mean theta = -0.05 (-0.0170, -0.0307, -0.1022) - no change to PCSInstrument.conf
      • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1
        • The files needed are
          • /home/telescope/TCS/Configuration/GCS/MODS_{L,R}.cfg
          • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
          • image files: /lbt/data/share/tcs/lbto.YYYMMDD
    • DONE MODS AGW Transform data - ~1" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky)
    • DONE Verify Tansform data
      • DONE Collect transform data for verification
      • DONE Measure Rot ZP with MODS MOS masks
      • DONE Dither Test M52 or M71
      • DONE Confirmation new transform is acceptable
    • DONE Spectrophotometric standard all modes - medium
      • DONE Grating
      • DONE Prism
    • non-sidereal guiding
      • DONE Imaging with dithers
      • Long-slit
        • TBD
    • Test Science OB
      • DONE longslit
      • imaging
        • DONE Test without syncoffset and large dither (+/-30")
          • got co-pointing warnings
        • Test with synoffset and large dither (+/-45")
    • Dithering Repeatability with Pointing Mask to check DET XY
      • DONE Before Transform
      • After Transform
    • Skyflats
      • dual g, dual r
    • DONE Blind offset tests
    • Pointing Model - 2hr
    • Test offsets for MODS1 for better alignment - requires test OT scripts (J. Power) low priority
      • OT set up is more complicated - this task could be tried later in restart
    • DONE GCS and IDL have opposite field correction (IT 8315) - Medium/High (1-1.5hrs)
      • Requires 1" or better
      • run IDL task field_collect (same as measuring field aberrations)
      • take LUCI data first to cross-check IDL vs GCS
  • LUCIs (high)
    • Rotator center
      • DONE L1
      • L2
    • AGW1 Alignment (LUCI-1) (1 night) Overview of AGW procedure
      • Files needed:
        • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1:
        • /home/telescope/TCS/Configuration/GCS/LUCI_L.cfg
        • image files: /lbt/data/share/tcs/lbto.YYYMMDD
        • oac@oac: /lbt/oacontrol/current/etc/oacontrol.conf
      • DONE verify SH spots illuminated (0.5 hr)
      • DONE verify guiding hotspot center (1 hr)
      • DONE rotator zeropoint (requires MOS mask) - (0.5 hr)
        • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
        • Mean Theta all masks = 0.064 deg - no update to PCSInstrument.conf
        • Longslit Method (backup for no masks)
          • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
        • LUCI-1 ZP mean theta =0.067 - (2/3 are 0.05 or less, one was 0.105) no change to PCSInstrument.conf
      • DONE Collect transform data - (2hrs)
        • Stone A3 list (updated to remove out of range) - 16 stars collected
      • DONE Process transform data
      • verify transform data (1 hr)
        • DONEdata collection
        • Rot ZP check
    • AGW2 Alignment (LUCI-2) (1 night) Overview of AGW procedure
      • Files needed:
        • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1:
        • /home/telescope/TCS/Configuration/GCS/LUCI_R.cfg
        • image files: /lbt/data/share/tcs/lbto.YYYMMDD
          oac@oac: /lbt/oacontrol/current/etc/oacontrol.conf
      • DONE verify SH spots illuminated (0.5 hr)
      • DONE verify guiding hotspot center (1 hr)
      • rotator zeropoint (requires MOS mask) - (0.5 hr)
        • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
        • Mean Theta N masks =
        • Longslit Method (backup method)
          • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
      • Collect transform data - (2hrs)
        • Process transform data
      • verify transform data (1 hr)
        • data collection
        • Rot ZP check
    • Imaging check
      • dither in field (0.5hr)
      • dither out of field (0.5hr)
        • DONE L1 - NGC 040 (cycled through 5 filters, K,H,J, H2, HeI)
        • L2
    • non-sidereal Imaging Check
      • dither in field (0.5hr)
      • dither out of field (0.5hr) -
        • DONE L1 - Comet 69P CG (cycled through 3 filters, K,H,J)
        • L2
    • DONE check longslit - high (1 hr)
      • DONE Science-like OB for G210
      • DONE Science-like OB for G200
    • DONE check longslit - NS
    • DONE M31 observation check
      • Center slit on bulge, verify GS can be found in crowded field.
      • Offset 20' to a sky frame and verify return can place slit on original position and find the original guidestar
      • This is a test for an upcoming INAF program.
      • Location: In V1.36 (and later) of the LUCI OT Restart script
    • Twilight flats
    • Pointing Model - High (2hrs)
      • PtStarrs ~10th mag 12-20 stars
      • Poor seeing/clouds OK
    • Photometric Standards
      • L1
        • UT Sep 06 - 2 sets/airmasses
      • L2
    • DONE Test of LUCI software while observing
      • work with software to check issues in realtime with RTD, and other LUCI software
      • high cadence observations
      • cube save
  • Mixed Mode Tests
    • LBC-Blue/LUCI-2 NS Mixed Mode verification (0.5hr) - dependent on LUCI-2 release
    • LBC-Red/MODS-1 (0.5-1hr)
      • DONE PA 90 both dither test
      • Orthogonal PAs
  • LBTI (4 x 0.5 night - need not be consecutve)
    • Night 1: Close AO loop on bright star (bin 1).
    • Night 1: Perform focus sequence in H band, confirm in L band.
    • Night 1: Test observing script for standard LMIRCam sequence.
    • Night 2: Take OSCO-like data set.
    • Night 2: Check out other bins.
    • Nights 3 & 4: Test and excercise Gopt & NCPA on bin 1 star, find stable configuration, develop efficient recovery procedures in case of issues
  • AO (2 x 0.5 night - need not be consecutve)
    • System checkout (0.5 night - any half)
      • Diffraction limited verification
        • Verify Bin 1
        • Verify Bin 4 - some elongation, hit R=18
        • Test Science scripts (5 hrs)
          • sideereal
          • Non-sidereal
        • OSCO (0.5hr)
      • ESM verification
        • Test science scripts (2 hr)
    • Science check (0.5 night) - 1st half
  • PEPSI (0.5 night -high)
    • DONE Verify PFU alignment
    • DONE Pointing Model - 2hr Medium/High
      • Remove retro-reflecting dichroic (to give full FOV)
      • Take images with PFU (Active mode) -not a requirement, images come "free"
      • Need ~28 stars (10th mag or so)
        • PtStars (IRAF) will generate stars based on LST
        • minimized long azimuth slews
      • Analysis with TPOINT - D. Thompson
      • Can do in poor seeing/clouds
    • DONE Training observations for staff - 1-2hrs per person - High
      • Dress rehearsal last two nights
  • Caveats
    • LBCS
      • Caveats:
    • MODS
      • Caveats:
    • LUCIs
      • Caveats:
    • LBTI
      • Caveats:
    • PEPSI
      • Caveats:

Rainy Night Tasks
  • We could make some more active TMS tests with LBC.



-- BarryRothberg - 15 Jul 2021

Support Files

MMS files:

Field MMS RA Dec MaskID Manufacture
StoneO mods.36.StoneO.mms 21 37 22.232 00 31 54.19 535231 2016
StoneA mods.37.StoneA.mms 00 56 20.127 -01 14 12.57 57172 2016
Gaia 00+25 mods.2.g00d25ep.mms 00 00 00.139 +25 00 17.090 531713 2021
Gaia 20+25 mods.3.g2025ep2.mms 20 00 00.910 +25 00 28.100 531956 2021
Gaia 04+25 mods.4.g04d25ep.mms 03 59 58.669 +25 00 40.050 532732 2021

LMS Files:

Field RA (J2000) Dec (J2000) MASK ID
StoneO (new) 21 37 31.791 00 31 55.010 903033
StoneP (new) 22 35 49.616 01 31 33.950 933322
StoneA (new) 00 54 02.050 -00 32 34.460 921934

Topic revision: r66 - 25 Sep 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback